OpenStudy (anonymous):

Solve for x 9(x – 2) = 5(2 + x)

OpenStudy (anonymous):

(9x^2 - 18x) = (10 +5x)

OpenStudy (texaschic101):

9(x - 2) = 5(2 + x) -- distribute through the parenthesis 9x - 18 = 10 + 5x -- subtract 5x from both sides 9x - 5x - 18 = 10 -- add 18 to both sides 9x - 5x = 10 + 18 -- combine like terms 4x = 28 -- divide both sides by 4 x = 7 check.. 9(7 - 2) = 5(2 + 7) 9(5) = 5(9) 45 = 45 (correct) x = 7 any questions ?

OpenStudy (anonymous):

It says solve for x I'm not sure if that's an x in front of the 9 or part of the problem.

OpenStudy (anonymous):

nope thanks man

OpenStudy (texaschic101):

:)